Consider the IVP y - 4y + 4y = 0, y = -2, y(0) = 1 a. Solve the IVP analytically b. Using step size 0.1, approximate y(0.5
Answer #1

– weve given D£ + y^4 44+48=0 with 960) = -2; y(0) = 1 (a) Analytically... Given De is homogeneous 2X 20 So auxillary ean m0.5 ho So for dy वर =Z= f1769, 7) Intl = Ynt hflinghiton) & d 2 = 42-44 = g (X,Y,Z) dr Lonto = Zrt. h J (Xh Inita) with Ylo)=Forn=t. Yg = 0.20799 :: 410.5) = 0.20749 Byeuler metlog 3 Aralytically 9(0-5)=l-3591909) EC), error = Actual value-approximaThe answer is in the pic. If you have any doubt please comment in the comment box. If this solution helps, upvote to encourage us. Thanks:)